# NestedSGX: Bootstrapping Trust to Enclaves within Confidential VMs

### v1

@@ Abstract.
This paper introduces NestedSGX, which leverages virtual machine privilege level (VMPL), a recent hardware feature available on AMD SEV-SNP to enable the creation of hardware enclaves within the guest VM. Similar to Intel SGX, NestedSGX considers the guest OS untrusted for loading potentially malicious code. It ensures that only trusted and measured code executed within the enclave can be remotely attested. To seamlessly protect existing applications, NestedSGX aims for compatibility with Intel SGX by simulating SGX leaf functions. We have also ported the SGX SDK to NestedSGX, enabling the use of existing SGX toolchains and applications in the system. Performance evaluations show that context switches in NestedSGX take about 35,000 – 37,000 cycles, approximately 2-3 times that of Intel SGX. NestedSGX incurs minimal overhead in most real-world applications, with an average overhead below 5% for most workloads and 23.0% for I/O intensive workloads.
@@ 1. Introduction
At the core of the confidential computing stack is the trusted execution environment (TEE), which isolates the code and data of a confidential workload from other code running on the system, even at the highest privilege levels. In particular, Intel SGX (sgx, [n. d.]) provides protection for the trusted components of an application, known as enclaves, allowing only measured code to run within the TEEs [^1] Intel SGXv2 (McKeen et al., 2016) enables dynamic changes to the page attribute and dynamic loading of code, but with strict controls in place.. On the other hand, VM-based TEEs, such as AMD SEV (Sev-Snp, 2020), Intel TDX (Intel, 2020b), and ARM CCA (arm, [n. d.]) offer a fully backward-compatible confidential virtual machine (CVMs), enabling the execution of existing applications without modifications. However, placing entire VMs within TEEs is generally considered less desirable: a VM image is far more than just a kernel and an application – it includes a large number of system services. Consequently, it remains uncertain whether it is more secure than running the software on premises or on existing cloud infrastructure. Therefore, an interesting question still remains open for VM-based TEEs:
@@
Design. In this paper, we address this open question through the design of our system, called NestedSGX. It provides a defense-in-depth approach to attesting applications (Figure 1). The trusted portion of an application (i.e., the enclave) is loaded and executed within an isolated environment, separate from the feature-rich OS, similar to the enclave abstraction of Intel SGX. This trusted portion is measured during the loading process and can be attested to a remote user. The foundation of trust for measurement and attestation lies in a small section of trusted code (i.e., the security monitor), operating at the highest privilege level within the VM, whose integrity is also measured and attested with the assistance of the hardware-based TEE. This approach empowers us to dynamically measure and attest the application code, enabling us to bypass the complexities associated with the guest OS.
@@
To facilitate the protection of existing applications, NestedSGX prioritizes compatibility with Intel SGX. This involves simulating SGX leaf functions to manage the entire life cycle of enclaves, including creation, memory management, teardown, and context switches. Furthermore, we have successfully ported the Intel SGX SDK (Intel, 2020a) and the Rust SGX SDK (Wang et al., 2019) to the NestedSGX framework, allowing for smooth integration of existing SGX toolchains and applications.
We have implemented NestedSGX on top of AMD’s Linux Secure VM Service Module (SVSM) framework (SVSM, 2022; AMD, 2023c), without necessitating any modifications of the host software, such as KVM. We conducted a series of benchmarking tests on commercial AMD SEV-SNP hardware and observed that the cost of world switches is about 2×2$\times$2 ×3×3$\times$3 × that of Intel SGX. Moreover, NestedSGX introduces negligible overhead in most real-world applications, with less than 5% overhead for memory and computation-intensive tasks and around 23.0% overhead for I/O-intensive tasks, while the worst-case overhead is 35.9% compared with running in native CVM.
Contributions. The contributions of this paper are summarized as follows.
@@
    Design. We present the design of NestedSGX, incorporating the VMPL feature within the AMD SEV-SNP architecture. NestedSGX addresses the challenge of establishing trust in the applications while minimizing reliance on the potentially compromised guest OS.
@@ 2.1. Intel SGX
Figure 2. The enclave’s life cycle in Intel SGX.
The SGX hardware enables user space applications to set aside private memory regions of code and data. The allocation of the private memory is managed through a set of privileged (ring-0) ENCLS leaf functions, while a set of unprivileged (ring-3) ENCLU leaf functions allow applications to enter and execute within these regions (Figure 2). Additionally, the hardware records meta-information for the enclave and its thread to effectively manage the enclave’s life cycle. In the event of interrupts or exceptions, the hardware triggers an Asynchronous Enclave Exits (AEX) event. During an AEX, the enclave’s context is saved and the application context is restored. Upon handling the interrupt or exception, the enclave context is restored when the execution resumes within the enclave. The specific details involved in these operations are omitted for brevity, and the readers may refer to the official manual (Intel, [n. d.]) for more information.
SGX attestation. During enclave creation, the hardware measures the integrity of the enclave code and data. This allows for remote attestation, enabling external entities to verify the trustworthiness of the enclave. In most cases, every component of enclave code can be measured and attested. The root of trust for attestation resides in the attestation key, which is exclusively accessible by the Intel signed quote enclave (QE).
@@ 2.2. SEV-SNP and VMPL
SEV-SNP, which was introduced in 2020 (Sev-Snp, 2020), is the third generation of SEV and provides enhanced security against malicious manipulation of page mappings by the host. One of its key features is the Reverse Map Table (RMP), a structure located in secure memory that maps system physical addresses (sPAs) to guest physical addresses (gPAs). The RMP serves as a metadata table managed by the AMD-SP and plays a critical role in tracking the ownership of each system physical page. It establishes a global one-to-one mapping between sPAs and gPAs, ensuring that a page cannot be simultaneously mapped into multiple guests or multiple times within a single guest.
Virtual Machine Privilege Level (VMPL) is an optional feature within the SEV-SNP architecture that allows a guest VM to divide its address space into four distinct levels. These levels serve as hardware-isolated abstraction layers for the VM, with VMPL0 representing the highest privilege level and VMPL3 representing the lowest. Each hardware context, known as a Virtual Machine Save Area (VMSA), is associated with a specific VMPL. Moreover, different memory pages assigned to a guest can have varying permissions based on the VMPL. VMPLs are utilized to implement additional page permission checks and are independent of other x86 security features.
@@
SEV-SNP attestation. The SEV-SNP CVM is initialized from an unencrypted initial image. This image contains the boot code necessary for the guest VM but does not include any confidential information. During the launch process, the hypervisor requests the AMD-SP to install this initial set of pages in the guest VM. The AMD-SP cryptographically measures the content of these pages, along with the associated metadata, to ensure an accurate measurement of the initial guest memory layout. Following that, the boot code triggers the bootstrapping of the OS image, which is also a part of the measured initial image. The attestation report, signed by the AMD-SP firmware, allows a third party, such as the guest owner, to verify the state of the CVM running on an authentic AMD platform.
@@ 2.3. Threat model
NestedSGX is built on top of SEV-SNP and follows a similar threat model, which considers the host software as a potential source of harm or compromise. We place our trust in the underlying SEV-SNP hardware to protect the CVM from direct observation or tampering by the host software or other CVMs.
Considering the substantial code base of the guest OS within the CVM, it is not considered entirely trustworthy. Therefore, the primary goal of NestedSGX is to strongly maintain the integrity and confidentiality of the enclave, effectively protecting it from potential security compromises originating from both the host software and the untrusted guest OS. These entities may collaborate in their attempts to breach the enclave’s security.
Furthermore, this paper does not cover scenarios in which the enclave intentionally reveals secrets or the security risks associated with unsafe code, such as memory safety vulnerabilities within the enclave or attacks originating from the interface between the enclave and the application (e.g., COIN (Khandaker et al., 2020) or IAGO (Checkoway and Shacham, 2013) attacks). However, it is important to highlight that NestedSGX ensures that the enclave cannot compromise the operations or integrity of the guest OS, even in the presence of vulnerabilities.
The security monitor operates in kernel mode at VMPL0 and is responsible for enforcing memory isolation between mutually distrustful components. Additionally, the security monitor serves as the core root of trust for measurement (CRTM). To establish a foundation of trust, it is assumed that the security monitor is fully trusted.
@@ 3.1. Overview
Figure 3. NestedSGX overview. The security monitor and the enclave operates in the kernel and user mode at VMPL0, while the guest OS and App operates in the kernel and user mode at VMPL0. The transitions between VMPL0 and VMPL1 occur via the untrusted host.
@@
As shown in Figure 3, NestedSGX encompasses the following components within the CVM. The application is divided into the trusted part (enclave) and the untrusted part (App). The guest OS (such as popular Linux distributions) offers essential features to applications, such as file systems, networking, device drivers, and language runtimes. The security monitor is the root of trust for measurement and attestation of the enclave. It is also responsible for enforcing security isolation within NestedSGX. To let the users maintain control over the code running inside the CVM and bootstrap trust to the enclave within a feature-rich guest OS, NestedSGX is designed to fulfill the following requirements.
Firstly, the system establishes isolated memory regions exclusively used by the security monitor, the enclaves and the guest OS respectively. This ensures that the enclave and the monitor is secure even in scenarios where the guest OS operates in kernel mode. Specifically, the entire gPA space is divided to two parts: a continuous and fixed region exclusively used by VMPL0, and the remaining portion used by VMPL1. The security monitor and the enclave are executed at VMPL0, while the guest OS and App operate at VMPL1. Additionally, our design guarantees that the enclave cannot access the memory of the guest OS, despite running at a higher VMPL. The details of the memory isolation scheme will be presented in Sec. 3.2.
Secondly, in NestedSGX, only authenticated code and data can be loaded and processed within the enclave. To achieve this, NestedSGX enforces a mechanism where the enclave can only be loaded by the trusted security monitor, which also performs the measurement of the enclave’s integrity. While the guest OS retains the freedom to load and execute any code within its own address space, these code segments are not affirmed by the security monitor and therefore will not be attested. The details of enclave measurement and attestation will be presented in Sec. 3.4.
@@
Lastly, NestedSGX ensures that any transitions between the trusted and untrusted components are mediated by the security monitor. This guarantees that sensitive information is properly sanitized and safeguarded. To leverage existing SGX toolchains, NestedSGX employs the SGX programming model to manage the enclave’s life cycle. The details for secure enclave state transitions will be presented in Sec. 3.3.

### v2

# The Road to Trust: Building Enclaves within Confidential VMs

@@ Abstract.
To address this issue, this paper introduces NestedSGX, a framework which leverages virtual machine privilege level (VMPL), a recent hardware feature available on AMD SEV-SNP to enable the creation of hardware enclaves within the guest VM. Similar to Intel SGX, NestedSGX considers the guest OS untrusted for loading potentially malicious code. It ensures that only trusted and measured code executed within the enclave can be remotely attested. To seamlessly protect existing applications, NestedSGX aims for compatibility with Intel SGX by simulating SGX leaf functions. We have also ported the SGX SDK and the Occlum library OS to NestedSGX, enabling the use of existing SGX toolchains and applications in the system. Performance evaluations show that context switches in NestedSGX take about 32,000 – 34,000 cycles, approximately 1.33×1.$3$1.54×1.54$\times$1.54 × higher than that of Intel SGX. NestedSGX incurs minimal overhead in most real-world applications, with an average overhead below 2% for computation and memory intensive workloads and below 15.68% for I/O intensive workloads.
copyright: noneconference: ACM Conference on Computer and Communications Security; Due 15 May 2019; London, TBDjournalyear: 2019
@@ 1. Introduction
At the core of the confidential computing stack is the trusted execution environment (TEE), which isolates the code and data of a confidential workload from other code running on the system, even at the highest privilege levels. In particular, Intel SGX (sgx, 2022) provides protection for the trusted components of an application, known as enclaves, allowing only measured code to run within the TEEs [^1] Intel SGXv2 (McKeen et al., 2016) enables dynamic changes to the page attribute and dynamic loading of code, but with strict controls in place.. On the other hand, VM-based TEEs, such as AMD SEV (Sev-Snp, 2020), Intel TDX (Intel, 2020b), and ARM CCA (arm, 2022) offer a fully backward-compatible confidential virtual machine (CVM), enabling the execution of existing applications without modifications. However, placing entire VMs within TEEs is generally considered less desirable: a VM image is far more than just a kernel and an application – it includes a large number of system services. Consequently, it remains uncertain whether it is more secure than running the software on premises or on existing cloud infrastructure. Therefore, an interesting question still remains open for VM-based TEEs:
@@
A seminal work called vSGX (Zhao et al., 2022) was developed to virtualize Intel SGX on AMD SEV, ensuring full binary compatibility and enabling the execution of unmodified SGX programs in CVMs. vSGX adopts a two-VM approach where one VM handles the untrusted application while the other hosts the enclave. However, this emulation of SGX leaf instructions incurs significant overhead, particularly for control flow transfers across the boundary of the two VMs (e.g., EENTER and EEXIT). For instance, an empty ECall operation on vSGX takes approximately 1.5ms, which is around 160×160$\times$160 × slower than on native Intel SGX. This slow context switch severely limits the service throughput achievable within the enclave. In practical scenarios, this limitation becomes a critical performance bottleneck. To illustrate, vSGX is 6×6$\times$6 × slower than the baseline on the cURL benchmark, and launching a 256 MB size vSGX enclave takes about 5 minutes.
@@
Design. We take a further step in this direction by introducing our system, called NestedSGX. NestedSGX follows a defense-in-depth approach to attesting applications (as depicted in Figure 1), where both the enclaves and the guest OS run within the same CVM. Specifically, the trusted portion of an application (i.e., the enclave) is loaded and executed within an isolated environment, separated from the feature-rich OS, similar to the enclave abstraction of Intel SGX. During the loading process, the trusted portion is measured and can be attested to a remote user. The core root of trust for measurement (CRTM) and root of trust for reporting (RTR) rely on a small segment of trusted code, known as the security monitor. This security monitor operates at the highest privilege level within the CVM, and its integrity is also measured and attested with the assistance of hardware-based TEE. This approach empowers us to dynamically measure and attest the application code, enabling us to bypass the complexities associated with the guest OS.
@@
To facilitate the protection of existing applications, NestedSGX prioritizes compatibility with Intel SGX. This involves simulating SGX leaf functions to manage the entire life cycle of enclaves, including creation, memory management, teardown, and context switches. Furthermore, we have successfully ported the Intel SGX SDK (Intel, 2020a), the Rust SGX SDK (Wang et al., 2019) as well as the Occlum library OS (Shen et al., 2020) to the NestedSGX framework, allowing for smooth integration of existing SGX toolchains and applications. In addition, we have implemented the HotCalls (Weisse et al., 2017) optimization in Occlum, which allows specific OCALLs to be handled asynchronously within the enclave without exiting it. This optimization has led to improved performance for applications running on Occlum without requiring any modifications to their source code. Unlike vSGX, we stress that full binary compatibility with SGX is a non-goal of our paper.
We have implemented NestedSGX on top of AMD’s Linux Secure VM Service Module (SVSM) framework (SVSM, 2022; AMD, 2023c), without necessitating any modifications of the host software, such as KVM. We conducted a series of benchmarking tests on commercial AMD SEV-SNP hardware and observed that the cost of world switches is about 1.33×1.$3$1.54×1.54$\times$1.54 × higher than Intel SGX, which is nearly two orders of magnitude faster than vSGX. Moreover, NestedSGX introduces negligible overhead in most real-world applications, with less than 2% overhead for memory and computation-intensive tasks and less than 15.68% overhead for I/O-intensive tasks. Generally, NestedSGX exhibits moderate costs that are comparable to other enclave systems, such as Intel SGX. Consequently, we view it as a promising solution for addressing the critical challenge of performing secure computations within CVMs.
Contributions. The contributions are summarized as follows.
@@
    Design. We present the design of NestedSGX, incorporating the VMPL feature within AMD SEV-SNP. NestedSGX addresses the challenge of establishing trust in the applications while minimizing reliance on the potentially compromised guest OS.
@@ 2.1. Intel SGX
The SGX hardware enables user space applications to set aside private memory regions of code and data. The allocation of the private memory is managed through a set of privileged (ring-0) ENCLS leaf functions, while a set of unprivileged (ring-3) ENCLU leaf functions allow applications to enter and execute within these regions. Additionally, the hardware records meta-information for the enclave and its thread to effectively manage the enclave’s life cycle. In the event of interrupts or exceptions, the hardware triggers an Asynchronous Enclave Exits (AEX) event. During an AEX, the enclave’s context is saved and the application context is restored. Upon handling the interrupt or exception, the enclave context is restored when the execution resumes within the enclave.
SGX attestation. During enclave creation, the hardware measures the integrity of the enclave code and data. This allows for remote attestation, enabling external entities to verify the trustworthiness of the enclave. In most cases, every component of enclave code can be measured and attested. The root of trust for reporting resides in the attestation key, which is exclusively accessible by the Intel signed quote enclave (QE).
@@ 2.2. SEV-SNP and VMPL
SEV-SNP, introduced in 2020 (Sev-Snp, 2020), is the third generation of SEV and provides enhanced security against malicious manipulation of page mappings by the host. One of its key features is the Reverse Map Table (RMP), a structure located in secure memory that maps system physical addresses (sPAs) to guest physical addresses (gPAs). The RMP serves as a metadata table managed by the AMD-SP and plays a critical role in tracking the ownership of each system physical page to prevent the host from writing to encrypted guest pages. It establishes a global one-to-one mapping between sPAs and gPAs, ensuring that a page cannot be simultaneously mapped into multiple guests or multiple times within a single guest.
Virtual Machine Privilege Level (VMPL) is an optional feature within the SEV-SNP architecture that allows a guest VM to divide its address space into four distinct levels. These levels serve as hardware-isolated abstraction layers for the CVM, with VMPL0 representing the highest privilege level and VMPL3 representing the lowest. Each hardware context, known as a Virtual Machine Save Area (VMSA), is associated with a specific VMPL. Moreover, different memory pages assigned to a guest can have varying permissions based on the VMPL. VMPLs are utilized to implement additional page permission checks and are independent of other x86 security features.
@@
SEV-SNP attestation. The SEV-SNP CVM is initialized from an unencrypted initial image. This image contains the boot code necessary for the CVM but does not include any confidential information. During the launch process, the hypervisor requests the AMD-SP to install this initial set of pages in the CVM. The AMD-SP cryptographically measures the content of these pages, along with the associated metadata, to ensure an accurate measurement of the initial guest memory layout. Following that, the boot code triggers the bootstrapping of the OS image, which is also a part of the measured initial image. The attestation report, signed by the AMD-SP firmware, allows a third party, such as the guest owner, to verify the state of the CVM running on an authentic AMD platform.
@@ 2.3. Threat model
NestedSGX is built on top of SEV-SNP and follows a similar threat model, which considers the host software as a potential source of harm or compromise. We place our trust in the underlying SEV-SNP hardware to protect the CVM from direct observation or tampering by the host software or other CVMs. Specifically, we do not consider the recent attacks on SEV-SNP, such as the CacheWrap attack (Zhang et al., 2024), the WeSee attack (Schlüter et al., 2024a) and the Heckler attack (Schlüter et al., 2024b).
Since the user can attest the load-time correctness of the CVM with remote attestation, we assume that the attacker initially only controls all software and hardware external to the CVM. However, considering the substantial code base of the guest OS within the CVM, it is not considered entirely trustworthy. Therefore, the primary goal of NestedSGX is to strongly maintain the integrity and confidentiality of the enclave, effectively protecting it from potential security compromises originating from both the host software and the untrusted guest OS. These entities may collaborate in their attempts to breach the enclave’s security.
Furthermore, this paper does not cover scenarios in which the enclave intentionally reveals secrets or the security risks associated with unsafe code, such as memory safety vulnerabilities within the enclave or attacks originating from the interface between the enclave and the application (e.g., COIN (Khandaker et al., 2020) or IAGO (Checkoway and Shacham, 2013) attacks). However, it is important to highlight that NestedSGX ensures that the enclave cannot compromise the operations or integrity of the guest OS, even in the presence of vulnerabilities. The security monitor operates in kernel mode at VMPL0 and is responsible for enforcing memory isolation between mutually distrustful components. Additionally, the security monitor serves as the core root of trust for measurement (CRTM). To establish a foundation of trust, it is assumed that the security monitor is fully trusted.
@@ 3.1. Overview
Figure 2. NestedSGX overview. The security monitor and the enclave operate in the kernel and user mode at VMPL0, while the guest OS and App operate in the kernel and user mode at VMPL1. The transitions between VMPL0 and VMPL1 occur via the untrusted host.
@@
As shown in Figure 2, NestedSGX encompasses the following components within the CVM. The application is divided into the trusted part (enclave) and the untrusted part (App). The security monitor is the root of trust for measurement and reporting of the enclave. It is also responsible for enforcing security isolation within NestedSGX. The guest OS (such as popular Linux distributions) offers essential features to applications, such as file systems, networking, device drivers, and language runtimes. To let the users maintain control over the code running inside the CVM and bootstrap trust to the enclave within a feature-rich guest OS, NestedSGX is designed to fulfill the following requirements.
@@
Firstly, NestedSGX ensures that any transitions between the trusted and untrusted components are mediated by the security monitor. This guarantees that sensitive information is properly sanitized and safeguarded. To leverage existing SGX toolchains, NestedSGX employs the SGX programming model to manage the enclave’s life cycle. The details for secure enclave state transitions will be presented in Sec. 3.2.
Secondly, the system establishes isolated memory regions exclusively used by the security monitor, the enclaves and the guest OS respectively. This ensures that the enclave and the security monitor are secure even in scenarios where the guest OS operates in kernel mode. Specifically, the entire gPA space is divided to two parts: a continuous and fixed region exclusively used by VMPL0, and the remaining portion used by VMPL1. The security monitor and the enclave are executed at VMPL0, while the guest OS and App operate at VMPL1. Additionally, our design guarantees that the enclave cannot access the memory of the guest OS, despite running at a higher VMPL. The details of the memory isolation scheme will be presented in Sec. 3.3.
Lastly, in NestedSGX, only authenticated code and data can be loaded and processed within the enclave. To achieve this, NestedSGX enforces a mechanism where the enclave can only be loaded by the trusted security monitor, which also performs the measurement of the enclave’s integrity. While the guest OS retains the freedom to load and execute any code within its own address space, these code segments are not affirmed by the security monitor and therefore will not be attested. The details of enclave measurement and attestation will be presented in Sec. 3.4.
Table 1. Comparison with competing approaches.
Unlimited enclaves Multi-threading Exception handling SGX ecosystem No changes to hypervisor Low-overhead context switches
vSGX (Zhao et al., 2022)
Veil (Ahmad et al., 2024)
NestedSGX
Comparison with competing approaches (Table 1). vSGX (Zhao et al., 2022) effectively virtualizes Intel SGX on AMD SEV, ensuring complete binary compatibility and the execution of unmodified SGX programs. vSGX employs two separated CVMs to accommodate the untrusted application and the enclave. In contrast, NestedSGX adopts a more intuitive design that utilizes different VMPLs for memory isolation. The NestedSGX design offers the following benefits.
Firstly, vSGX cannot support directly sharing encrypted memory between the enclave VM (EVM) and application VM (AVM). To provide confidentiality, integrity and replay protection, a dedicated communication protocol and encrypted channel are essential. Within NestedSGX, the security monitor can access the entire guest physical address space, and communicate with the guest OS through shared memory. This approach circumvents the overhead associated with memory encryption and movement, and prevents the hypervisor from observing traffic patterns. As a comparison, an empty ECALL costs 1.5 ms on vSGX, and only 12 us on NestedSGX, which is two orders of magnitude faster than vSGX.
Secondly, with vSGX, only one enclave can be hosted within the EVM, which restricts the maximum number of concurrently-running enclaves on the hardware platform. This limitation arises because SEV associates ASIDs with VMs’ memory encryption keys, and the ASID bit is limited (4 bits (Gu et al., 2020)). In contrast, NestedSGX has the ability to support an unlimited number of enclaves.
More importantly, NestedSGX does not serve as a direct substitute for vSGX. On vSGX, the enclave thread and application thread operate with different vCPUs in separated VMs. Specifically, the application thread at the AVM is paused on EENTER; upon receiving the EENTER request, the EVM creates a thread to handle it in the EVM. On NestedSGX, the enclave thread and application thread operate with the same vCPU, while the vCPU contexts associated with different VMPLs are saved in different VMSAs. Consequently, managing the entry and exit of enclaves on NestedSGX necessitates meticulous handling of vCPU state transitions. As an evidence, AEX is relatively straightforward on vSGX: vSGX simply puts the enclave thread to sleep on AEX, then wakes it and resumes the execution directly on ERESUME. On NestedSGX the enclave context needs to be saved on AEX and recovered on ERESUME.
Concurrent to our work, Ahmad et al. propose Veil (Ahmad et al., 2024) (to be presented at ASPLOS 2024) as a service framework running at VMPL0. The concept of Veil is quite similar to AMD’s Secure VM Service Module (SVSM). As one of the use cases, Veil offers protection to the entire applications within a separated VMPL, facilitating redirection of system calls and interrupts. However, unlike our approach, it does not focus on protecting a trusted portion of the application or strive for compatibility with Intel SGX. Currently, Veil lacks support for multi-threading and exception handling. In contrast, NestedSGX leverages an existing SGX toolchain by emulating most SGX instructions, thereby naturally providing support for multi-threading and exception handling.
3.2. Enclave life cycle management
In NestedSGX, the life cycle of an enclave closely resembles that of SGX. Initially, the application invokes the ECREATE leaf function to create the enclave, which initializes the SGX Enclave Control Structure (SECS) page. Subsequently, EPC pages necessary for the enclave, such as code sections, data sections, and Thread Control Structure (TCS) pages, are added using EADD and measured using EEXTEND leaf functions. Once all the required EPC pages are added, the enclave is initialized with EINIT, and the final measurement is performed. Upon completion of the enclave’s execution, the EPC pages are reclaimed using the EREMOVE leaf function, and the enclave is destroyed.
NestedSGX achieves compatibility with Intel SGX by transitioning to the security monitor on these leaf functions. Since VMPL switches can only be executed in the privileged mode (Sec. 2.2), NestedSGX provides the NestedSGX-driver within the guest OS, which accepts the request from the App and facilitates the VMPL switch. During this process, the NestedSGX-driver receives the type of SGX leaf functions, while the parameters are saved on the App’s stack. The NestedSGX-driver then switches the execution to VMPL0 and transfers control to the security monitor, where the security monitor emulates the instructions strictly adhering to the SGX specifications.
Synchronous enclave entry and exit. Once an enclave is initialized, the App can enter the enclave using the EENTER leaf function and jump to the enclave’s code for execution. As shown in Figure  3, during the emulation of EENTER, the security monitor switches to the enclave’s page table and enters the enclave entry point specified by the TCS using the sysret instruction. The enclave can then return to the App using the EEXIT leaf function, which is also emulated by the security monitor. Specifically, the enclave returns to the security monitor using the syscall instruction. Subsequently, the security monitor switches the execution to the NestedSGX-driver running at VMPL1. The NestedSGX-driver restores the App’s page table and returns control to the App using the sysret instruction.
Asynchronous enclave entry and exit. During enclave execution, faults and exceptions may occur. In accordance with the SGX model, these faults and exceptions are supposed to be handled by the guest OS first. In NestedSGX, when a fault or exception occurs within the enclave, it traps to the fault and exception handler within the security monitor (Figure 3). The security monitor then emulates the Asynchronous Exit (AEX) event by saving the context of the enclave in the state saving area (SSA). Subsequently, the security monitor switches the vCPU to the NestedSGX-driver, which operates at VMPL1. The NestedSGX-driver fills the instruction pointer (RIP) to the trampoline area (AEP) and invokes the handler of the guest OS. After the OS handles the fault or exception, it uses the iret instruction to return to the AEP. The AEP, in turn, resumes the execution of the enclave using the ERESUME leaf function. The emulation of the ERESUME leaf function is similar to that of EENTER, except that the enclave’s execution context is restored from the SSA. As a result, NestedSGX is capable of handling exceptions that occur during enclave execution. For example, NestedSGX supports the emulation of the CPUID instruction, enabling the execution of unmodified applications. [^2] In comparison, vSGX (Zhao et al., 2022) requires modifying the application code in order to bypass the CPUID check.
It’s worth noting that if the untrusted NestedSGX-driver chooses not to fill RIP with AEP, this decision does not introduce any new security issues. This is because the control flow outside the enclave is already vulnerable to manipulation within the SGX model and is not inherently trusted. In NestedSGX, faults and exceptions are firstly trapped to the security monitor, enabling easy detection of suspicious behaviors such as abnormal interrupts caused by page fault-based attacks. Conceptually, NestedSGX can also support the recent AEX-notify feature (Constable et al., 2023), although the implementation is planned for future work.